OpenStudy (anonymous):

6. An electric iron has a rating of 750W,220V. Calculate : a) current passing through it and b) its resistance when in use. 7. Two lamps , one rated 100W at 220V and other 40 W at 220V are connected parallel to 220V mains supply. Calculate the current drawn from the supply line. 8. An electric lamp is marked 100W,220V.It is used for 5 hours daily.Calculate: a)Its resistance when in use. b)Energy consumed in KWh per day. 9.two resistors of resistance 3 ohm and 6 ohm respectively are connected to a battery of 6V so as to have a) Min resistance

OpenStudy (kropot72):

6. a) Power (watts) P = V * I I = P/V b) \[P=I ^{2}R\] \[R=\frac{P}{I ^{2}}\] 7. Calculate current for ewach lamp using I = P/V and add the two values of current to find the total current. 8. a) Current I = P/V \[R=\frac{P}{I ^{2}}\] b) 100W = 0.1kW Energy = 0.1 * 5 kWh per day 9. Connecting the resistors in parallel gives minimum resistance.

OpenStudy (kropot72):

* for each lamp

OpenStudy (anonymous):

thank you for your guidance kropot72.

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

could you explain Q9. again, please?

OpenStudy (kropot72):

Q9 is not the full question? Please check.

OpenStudy (anonymous):

actually, I had reached the limit of the page, so I could type no more.

OpenStudy (kropot72):

You can post it as a new question.

OpenStudy (anonymous):

yes, I will.Thank you.